Removed ground from DSSD #1 to the chamber and set MACBs so that timestamp is from the NIM WR emulator not the VME VETAR2.
Root MACB mode 0x3 -> 0xd - all other MACB modes 0x3 ( unchanged ).
No timestamps observed from the emulator. Re-set to take timestamps from the VETAR" unit -> all timestamps observed.
10:30
No pulser observed in nn FEE modules 9, 10, 13, 14.
ADC data items check OK - all ~3k - attachment 2
Rates - all OK - attachment 3
1.8L spectra. - attachment 4,5
1.8W waveforms - attachment 6,7
11:00
Checked daisy chain which goes from 8->14->13->10->9. Replaced the lemo cable from 8->14 and pulsers re-appear. Rates all OK except aida10 which has now disappeared, tried asic check and reset and startup again no change in aida10.
Now majority of pulsers are observable the widths will be noted in a spreadsheet
After replacing lemo for nn pulser chain, all nn pulser widths decreased slightly ~10% for all nn FEEs, little change in pn pulser widths but n pulser for DSSD#1,2 appears more gaussian
11:44
FEE64s power cycled. aida10 now prioducing data.
Removed internal jumper from CAEN N1419ET
ADC data items check OK - all ~3k - attachment 8
Rates - all OK - attachment 9
1.8L spectra - attachment 10,11 little change in pn pulser widths but nn pulser widths are slightly broader.
1.8W waveforms - attachment 12,13
15:50 Jumper re-mounted onto caen bias module
ADC data items check OK - all ~3k - attachment 14
Rates - all OK - attachment 15
1.8L spectra - attachment 16,17
1.8W waveforms - attachment 18,19
16:45 Ground cables attached from water pipe (chamber ground) to lemo cable on aida08 (nn), aida13 (nn), aida09 (nn), aida01 (pn) so that all DSSDs are connected to chamber ground.
ADC data items check OK - all ~3k - attachment 20
Rates - all OK - attachment 21
1.8L spectra - attachment 22,23
1.8W waveforms - attachment 24,25
|